# Description length of canonical and microcanonical models

### v2

@@ B.1 Bayesian-Jeffreys description lengths
$\text{DL}^{\text{B, Jeff}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\frac{1}{2}\ln\frac{\pi nm}{2}+O\left(1/nm\right).$ (106)
The last expression is compared with the asymptotic expansion of the canonical NML description length $\text{DL}^{\text{NML}}_{\text{can}}$, obtained as the difference between the asymptotic expansion of (94) and the canonical log-likelihood:
$\text{DL}^{\text{NML}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\frac{1}{2}\ln\frac{\pi nm}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi nm}}+o(1).$ (107)
Finally, Equation (82) is obtained as the difference between (107) and (106)
B.2 Proving identities
@@
$\text{DL}_{\text{can}}^{\text{NML}}=\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}},$ (108)
@@
$\hat{W}(\mathbf{c})$ $=\frac{\sum_{\mathbf{G}:\mathbf{c}(\mathbf{G})=\mathbf{c}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G})}{\sum_{\mathbf{G}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G}))}$ (109)
@@
$\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}}(\mathbf{G}^{*})$ $=\ln\Omega(\mathbf{c}^{*})-\ln\hat{W}(\mathbf{c}^{*})=$ (110)
@@
$\text{DL}_{\text{mic}}^{\text{NML}}=\text{DL}_{\text{can}}^{\text{Bayes-Uniform}},$ (111)
@@
$w^{\text{Uniform}}(p)=1\quad\text{for }p\in[0,1].$ (112)
@@
$w^{\text{Uniform}}(\mathbf{p})=1\quad\text{for }\mathbf{p}\in[0,1]^{n}$ (114)
@@
$=\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*}),$ (116)

### v3

@@ Acknowledgements
$\text{DL}^{\text{Bayes-Jeffreys}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\frac{1}{2}\ln\frac{\pi nm}{2}+O\left(1/nm\right).$ (S20)
The last expression is compared with the asymptotic expansion of the canonical NML description length $\text{DL}^{\text{NML}}_{\text{can}}$, obtained as the difference between the asymptotic expansion of (S8) and the canonical log-likelihood:
$\text{DL}^{\text{NML}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\frac{1}{2}\ln\frac{\pi nm}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi nm}}+o(1).$ (S21)
Finally, Equation (82) is obtained as the difference between (S21) and (S20)
S2.2 Proving identities
@@
$\text{DL}_{\text{can}}^{\text{NML}}=\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}},$ (S22)
@@
$\hat{W}(\mathbf{c})$ $=\frac{\sum_{\mathbf{G}:\mathbf{c}(\mathbf{G})=\mathbf{c}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G})}{\sum_{\mathbf{G}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G}))}$ (S23)
@@
$\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}}(\mathbf{G}^{*})$ $=\ln\Omega(\mathbf{c}^{*})-\ln\hat{W}(\mathbf{c}^{*})=$ (S24)
@@
$\text{DL}_{\text{mic}}^{\text{NML}}=\text{DL}_{\text{can}}^{\text{Bayes-Uniform}},$ (S25)
@@
$w^{\text{Uniform}}(p)=1\quad\text{for }p\in[0,1].$ (S26)
@@
$w^{\text{Uniform}}(\mathbf{p})=1\quad\text{for }\mathbf{p}\in[0,1]^{n}$ (S28)
@@
$=\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*}),$ (S30)